OpenStudy (anonymous):

The sum of the integers p and q is 495. If p is divided by 10, and the result is equal to q, then what is p?

OpenStudy (abb0t):

p + q = 495 (sum means to add) \[\frac{ p }{ 10 }\] setting THAT equal to q. thus: \[p + \frac{ p }{ 10 } = 495\]

OpenStudy (anonymous):

p+q=495 p/10=q p=10q plug that in first equation 1oq +q=495 11q=495 q=45 p=450

OpenStudy (abb0t):

I forgot to mention that \[q = \frac{ p }{ 10 }\] which is why you substituted that for q in the original equation.

OpenStudy (anonymous):

and solve for p?

OpenStudy (abb0t):

Yes. Also, don't forget to solve for "P". Correct :)

OpenStudy (anonymous):

@abb0t thnx bro